# Gravitational-wave signals for supernova explosions of three-dimensional progenitors

### v1

@@ III.2 Signal in the frequency domain
Above the prominent f/g-mode strip, one can recognize weaker, broad-band emission extending to very high frequencies $f\gtrsim 3500\,\operatorname{Hz}$ in s12.28. This high-frequency “haze” is caused by a superposition of different PNS oscillation modes, whose precise nature is still unclear [66]. Rotation in the PNS accretion layer due to the large angular momentum gained from accretion downflows (see Figs. 24 and discussion in Sec. II) may also play a role for the contributing oscillation modes and the exact structure of the GW emission spectrum at very high frequencies. The spectrogram for model s12.28 exhibits a fainter, narrow band at $f\sim 1100-1300\,\operatorname{Hz}$ between roughly 0.3 s and 0.7 s. This gap-like feature is interpreted as a consequence of an avoided crossing of PNS f- and g-mode pulsations, as mentioned above already and extensively discussed before in Refs. [51, 54, 66].
@@ IV.1 Signal in the time domain
One can witness variability of the anisotropy parameters in Fig. 11 on different time scales: high-frequency fluctuations with ${\cal O}(10\,\mathrm{ms})$ periods as well as longer-time-scale modulations with lengths of hundreds of ms to seconds. The neutrino GW strains and amplitudes of Eqs. (12) and (15), however, give rise to a low-frequency neutrino GW memory that is imprinted on spacetime by the anisotropically expanding flux of emitted neutrinos [10, 11, 77]. Because of the time integration performed in Eq. (12), excursions in $\alpha_{\sigma}$ with opposite signs cancel each other out and fluctuations with the same sign sum up to monotonic trends.
@@ V.1 Comparison of matter GW signals
We further note that we do not witness any systematic absence of oscillations in $A_{\times}$ during the prompt postshock convection phase at $t_{\mathrm{pb}}\lesssim 50$ ms as reported for the Fornax models in Ref. [66]. In our simulations we find that $A_{+}$ can be weak in some viewing directions, but usually the amplitudes for both polarizations are present with similar strengths (see Fig. 6). Since the 3D nature of the progenitors in our simulations has an impact only when the convective oxygen-burning shell falls through the SN shock (i.e., typically at $t_{\mathrm{pb}}\gtrsim 0.2$ s), we wonder whether this difference could be connected to the different numerical grids used, namely a dendritic grid with mesh coarsening in the Fornax models versus a Yin-Yang grid without any preferred axis directions in our computations. Also larger or smaller grid-induced numerical perturbations might play a role in explaining the spread of magnitudes found for the GW amplitudes shortly after core bounce with the Fornax, CoCoNuT-FMT, Chimera, and our Prometheus-Vertex codes.
@@
Another difference concerns the onset of the explosions in correlation with the vigor of the GW emission. In the Fornax models the onset of the explosion (defined by the time when the stalled shock begins to expand outward) in all cases except the lowest-mass (9 $M_{\odot}$) models takes place during a phase of relatively weak matter-associated GW activity before the GW emission becomes significantly more powerful (see Figs. 3 and 4 in [66]). The corresponding moments in our simulations are $t_{\mathrm{pb}}\sim 170$ ms for s12.28 and $t_{\mathrm{pb}}\sim 250$ ms for s18.88, see Fig. 1.888It is important to note that we usually define —following the convention in previous publications— the onset of the explosion, $t_{\mathrm{exp}}$, as the time after bounce when the angle-averaged shock radius exceeds 400 km (which is roughly the time of no shock return). This is the reason why the times given here differ from those listed for $t_{\mathrm{exp}}$ in Table 1. In contrast to the Fornax models, these instants are within the time intervals of strong GW emission associated with non-radial hydrodynamic mass motions in the neutrino-heated postshock volume. (In this respect our models resemble the non-rotating 15 $M_{\odot}$ models in [30].) While the Fornax models show a substantial growth of the GW amplitudes during the PNS accretion that follows the onset of the explosions, our models yield at most a slight increase of the GW amplitudes after the SN shock has taken off (Figs. 6 and 7). These differences could be indicative of considerably stronger neutrino heating in the simulations of Refs. [66, 67].

### v2

@@ III.2 Signal in the frequency domain
Above the prominent f/g-mode strip, one can recognize weaker, broad-band emission extending to very high frequencies $f\gtrsim 3500\,\operatorname{Hz}$ in s12.28. This high-frequency “haze” is caused by a superposition of different PNS oscillation modes, whose precise nature is still unclear [66]. Rotation in the PNS accretion layer due to the large angular momentum gained from accretion downflows (see Figs. 24 and discussion in Sec. II) may also play a role for the contributing oscillation modes and the exact structure of the GW emission spectrum at very high frequencies. The spectrogram for model s12.28 exhibits a fainter, narrow band at $f\sim 1100-1300\,\operatorname{Hz}$ between roughly 0.3 s and 0.7 s. This gap-like feature is interpreted as a consequence of f- and p-mode PNS pulsations interfering with g-mode oscillations, as extensively discussed in Refs. [51, 54, 123, 61, 66].
@@ IV.1 Signal in the time domain
One can witness variability of the anisotropy parameters in Fig. 11 on different time scales: high-frequency fluctuations with ${\cal O}(10\,\mathrm{ms})$ periods as well as longer-time-scale modulations with lengths of hundreds of ms to seconds. [^6] Interestingly, we do not witness any clear correlations between features in the neutrino anisotropy parameters $\alpha_{+}$ and $\alpha_{\times}$ (Fig. 11) and the phases of enhanced PNS accretion (Fig. 5), which increase the angular momentum in layers around the neutrinopheres. In contrast, in an analysis of PNS kicks in Ref. [95], we found a strong correlation between phases of higher PNS accretion and excursions of the instantaneous neutrino dipole-asymmetry parameter $\alpha_{\nu}^{\mathrm{tot}}$ that is relevant for the neutrino-induced PNS kick (see Figs. 11 and 17 for model s12.28 in [95]). We understand this discrepancy by the fact that at late times the accretion geometry is usually dominated by a dipolar configuration with a massive downflow in one hemisphere and an outflow in the other hemisphere (see Fig. 4). Since in the late evolution phase the asymmetry of the neutrino radiation seen at large distances is mostly caused by neutrino absorption in the dense downflows rather than asymmetric neutrino emission from the neutrinospheres (as discussed in [95]), the accretion asymmetry has a substantial impact on the neutrino dipole parameter $\alpha_{\nu}^{\mathrm{tot}}$ but not prominently on the neutrino anisotropy parameters $\alpha_{\sigma}$ for the neutrino GW production. The neutrino GW strains and amplitudes of Eqs. (12) and (15), however, give rise to a low-frequency neutrino GW memory that is imprinted on spacetime by the anisotropically expanding flux of emitted neutrinos [10, 11, 77]. Because of the time integration performed in Eq. (12), excursions in $\alpha_{\sigma}$ with opposite signs cancel each other out and fluctuations with the same sign sum up to monotonic trends.
@@ V.1 Comparison of matter GW signals
We further note that we do not witness any systematic absence of oscillations in $A_{\times}$ during the prompt postshock convection phase during the first $\sim$ 50 ms after bounce as reported for the Fornax models in Ref. [66]. In our simulations we find that $A_{+}$ can be weak in some viewing directions, but usually the amplitudes for both polarizations are present with similar strengths (see Fig. 6). Since the 3D nature of the progenitors in our simulations has an impact only when the convective oxygen-burning shell falls through the SN shock (i.e., typically at $t_{\mathrm{pb}}\gtrsim 0.2$ s), we wonder whether this difference could be connected to the different numerical grids used, namely a dendritic grid with mesh coarsening in the Fornax models versus a Yin-Yang grid without any preferred axis directions in our computations. Also larger or smaller grid-induced numerical perturbations might play a role in explaining the spread of magnitudes found for the GW amplitudes shortly after core bounce with the Fornax, CoCoNuT-FMT [30], Chimera [57], and our Prometheus-Vertex codes.
The matter GW signals attributed to prompt postshock convection in Ref. [66] also stick out by their timing in the Fornax models, where in all cases this emission starts at several 10 ms after bounce, in some cases even at $t_{\mathrm{pb}}\gtrsim 50$ ms. This is much later than in our models and in those of Refs. [30, 57]. In fact, these times are too late for prompt convection behind the decelerating core-bounce shock, because the shock stalls within only milliseconds after bounce, thus forming a negative entropy gradient in its downstream region that immediately triggers convective overturn. Typically this early convective activity and its GW production have ended after 30–50 ms, as visible in Fig. 6. Instead, the powerful GW activity at several 10 ms post bounce in the Fornax models might originate from grid-induced numerical perturbations or it might indicate a strong transient associated with hydrodynamic relaxation that could be connected to a mapping from 1D collapse simulations to the 3D post-bounce modeling at too late times after core bounce. Accordingly, these earliest phases of vigorous GW emission are followed by only short periods of relative GW quiescence before enhanced GW emission develops again due to the hydrodynamic instabilities that occur in the neutrino-heated postshock layer.
@@
Another difference, possibly linked to the phenomenon described above, concerns the start of the explosions in correlation with the vigor of the GW emission. In many of the Fornax models —except the lowest-mass ($\sim$ 9 $M_{\odot}$) cases— the explosion sets in (defined by the time when the stalled shock begins to expand outward) during a phase of relatively weak matter-associated GW activity. This instant is immediately or soon after the earliest phase of strong GW production, but still before the GW emission regains more power by neutrino-driven postshock convection (see Figs. 3 and 4 in [66]). The corresponding moments of the onset of explosions in our simulations are $t_{\mathrm{pb}}\sim 170$ ms for s12.28 and $t_{\mathrm{pb}}\sim 250$ ms for s18.88, see Fig. 1.999It is important to note that we usually define —following the convention in previous publications— the onset of the explosion, $t_{\mathrm{exp}}$, as the time after bounce when the angle-averaged shock radius exceeds 400 km (which is roughly the time of no shock return). This is the reason why the times given here differ from those listed for $t_{\mathrm{exp}}$ in Table 1. In contrast to the Fornax models, these points are within the time intervals of strong GW emission associated with non-radial hydrodynamic mass motions in the neutrino-heated postshock volume. (In this respect our models resemble the non-rotating 15 $M_{\odot}$ models in [30].) While the Fornax models show a substantial growth of the GW amplitudes during the PNS accretion that follows the onset of the explosions, our models yield at most a slight increase of the GW amplitudes after the SN shock has taken off (Figs. 6 and 7). These differences could be indicative of considerably stronger neutrino heating in the simulations of Refs. [66, 67].